OpenStudy (anonymous):

Which expressions are equal to the expression 5.3 – 14.2? Choose all answers that are correct. A. 5.3 + (–14.2) B. –5.3 + 14.2 C. –14.2 + (–5.3) D. –14.2 – (–5.3)

OpenStudy (anonymous):

Hello! To solve this you need to solve the problem and then the answer choices. I can do the first one for you 5.3-14.2=-8.9 A. 5.3+(-14.2)= -8.9 Woop there's your answer!

OpenStudy (anonymous):

Oh!! WELCOME TO OPEN STUDY! CONGRATS ON THE FIRST QUESTION! YAY!

OpenStudy (anonymous):

Thanks and hello and its it only a because it can be more than one :D

OpenStudy (anonymous):

Oh it's one of those multiple choice-multiple choice answer things. Okay so lets just solve them both together B. -5.2+14.2=-9 C. -14.2+(-5.3)=-19.5 D. -14.2-(-5.3)=-8.9 :D

OpenStudy (anonymous):

So its be D and A